OpenStudy (amina_soneviseth):

MEDAL AND FAN 6 ( x + 1/3 ) = 6 x + 2/3

OpenStudy (anonymous):

best bet is to distribute the \(x\) on the left, get \[6x+2=6x+\frac{2}{3}\] as a start

OpenStudy (anonymous):

oops i meant "distribute the 6"

OpenStudy (anonymous):

then do not go any further you have \(6x\) on the left and on the right whatever \(x\) is it is the same on both sides and theerfore so is \(6x\) you cannot add \(2\) to a number and get the same answer if you add \(\frac{2}{3}\) to it, so you cannot solve this

OpenStudy (amina_soneviseth):

okay so thats 6x + 2 = 9? :D then what @satellite73

OpenStudy (anonymous):

unless it was \[6(x+\frac{1}{2}=\frac{6x+2}{3}\] which is a different story

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

is the question \[6(x+\frac{1}{2})=6x+\frac{2}{3}\] like that?

OpenStudy (amina_soneviseth):

yes? i kind of did the problem... its zero solution! thank you for the help! :D

OpenStudy (anonymous):

NO solution yes

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

love your ipod cases btw

OpenStudy (amina_soneviseth):

?? what did you mean? @satellite73

OpenStudy (amina_soneviseth):

ohh! haha thank you! how did you fins that? :D @satellite73

OpenStudy (amina_soneviseth):

*find